Cabio Company manufactures two products, Product C and Product D. The company estimated it would incur $181,910 in manufacturing overhead costs during the current period. Overhead currently is applied to the products on the basis of direct labor-hours. Data concerning the current period’s operations appear below:

Product C Product D
Estimated volume 4,000 units 3,200 units
Direct labor-hours per unit 1.60 hours 1.70 hours
Direct materials cost per unit $ 13.60 $ 26.10
Direct labor cost per unit $ 16.00 $ 17.00

Required:

a-1. Compute the predetermined overhead rate under the current method. (Round your answer to 2 decimal places.)

a-2. Determine the unit product cost of each product for the current year. (Round your intermediate calculations and final answers to 2 decimal places.)
Answer #1

1) $15.36 per hour

predetermined overhead rate=estimated manufacturing overhead/estimated activity level

estimated manufacturing overhead is $181910

estimated activity level total direct labor hours

direct labour hours=Estimated volume*Direct labor-hours per unit

for

product c=4000*1.60=6400 hours

product d=3200*1.70=5440 hours.

total direct labour hours=6400+5440=11840 hours

predetermined overhead rate=$181910/11840=$15.36 per hour

2)

unit product cost=direct material cost per unit + Direct labor cost per unit + manufacturing overhead per unit.

manufacturing overhead per unit=predetermined overhead rate*Direct labor-hours per unit

for

product c=$15.36*1.6=24.57

product d=$15.36*1.7=26.11

unit product cost for

product c=$13.60+$16+$24.57=$54.17 per unit

product d=$26.10+$17+26.11=$69.21 per unit